P= $3,000
R = 6%
T = 3 years
What is the interest you would make?

Answer by rfer(16322)

Simple interest
I=Prt
I=3000*0.06*3
I=$540
Compounded
A=3000(1.06)^3
A=3000*1.191
A=$3573
3573-3000=$573
You didn't say which